Ans please with Solution..
A farmer uses 140 feet of fencing to enclose a rectangular field, If the length of the field is (33)1/3% greater than its width .What is the diagonal distance from one corner of the enclosure to the opposite corner?

Answer by rothauserc(4718)

33 and (1/3)% = 33 and (1/3) / 100 = (33 and 1/3 / 100) * (3/3) = 100 / 300 = 1/3
:
length = width(w) + (w/3)
:
perimeter(P) = 2w + 2(4w/3)
:
140 = 2w + 8w/3
:
14w/3 = 140
:
14w = 560
:
w = 40 and l = 160/3
:
use Pythagorean Theorem to find the length of the diagonal
:
(120/3)^2 + (160/3)^2 = d^2
:
d^2 = 40000 / 9
:
d = square root (40000 / 9) = 200 / 3 = 66 and (2/3) feet
:
**************************************
the diagonal is 66 and (2/3) feet long
***************************************